What volume of NH3 gas in liters is produced at STP when 0.61 L of a 1.81 M NH4Cl solution reacts according to the above equation?


1
Expert's answer
2021-04-12T06:28:14-0400

NH4Cl = NH3 + HCl

Mole = volume × concentration

= 0.61L × 1.81M = 1.1041mole

1 mole NH4Cl = 22.4 L NH3

1.1041 mole NH4Cl =1.1041 × 22.4L NH3

= 24.73L NH3